# QPET: A Versatile and Portable Quantity-of-Interest-Preservation Framework for Error-Bounded Lossy Compression

### v3

@@ Abstract.
Error-bounded lossy compression has been widely adopted in many scientific domains because it can address the challenges in storing, transferring, and analyzing unprecedented amounts of scientific data. Although error-bounded lossy compression offers general data distortion control by enforcing strict error bounds on raw data, it may fail to meet the quality requirements on the results of downstream analysis, a.k.a. Quantities of Interest (QoIs), derived from raw data. This may lead to uncertainties and even misinterpretations in scientific discoveries, significantly limiting the use of lossy compression in practice. In this paper, we propose QPET, a novel, versatile, and portable framework for QoI-preserving error-bounded lossy compression, which overcomes the challenges of modeling diverse QoIs by leveraging numerical strategies. QPET features (1) high portability to multiple existing lossy compressors, (2) versatile preservation to most differentiable univariate and multivariate QoIs, and (3) significant compression improvements in QoI-preservation tasks. Experiments with six real-world datasets demonstrate that integrating QPET into state-of-the-art error-bounded lossy compressors can gain 2x to 10x compression speedups of existing QoI-preserving error-bounded lossy compression solutions, up to 1000% compression ratio improvements to general-purpose compressors, and up to 133% compression ratio improvements to existing QoI-integrated scientific compressors.
copyright: none
@@ 1. Introduction
Today’s high-performance computing facilities and high-resolution instruments produce vast amounts of scientific data that overwhelm the data storage and transmission systems. According to recent studies, climate simulations generate hundreds of TB of data every 16 seconds (Foster et al., 2017), and fusion applications may produce 1 EB of data in a single run with current- and next-generation exascale computing systems (fus, 2021). In the molecular dynamics (MD) domain, the EXAALT project produces trajectory data with a trillion time steps (exa, 2020), which require hundreds of terabytes of disk space in a scientific data management system (e.g., HDF5 (hdf, [n.d.])). Those examples pose grand challenges in the underlying data management tasks, including I/O and data transfer, which necessitates the need for effective data compression.
@@
Data compression has been widely used in reducing data volumes (Wiegand et al., 2003; Sze et al., 2014; Pelkonen et al., 2015; Jensen et al., 2018) and accelerating queries (Chen et al., 2001; Arion et al., 2007; Zhang et al., 2018, 2021) in database and data management systems. Nonetheless, challenges arise when applying existing compression techniques in the context of scientific data. On the one hand, while lossless compression techniques (Pelkonen et al., 2015; Alakuijala et al., 2018; Lindstrom et al., 2017; Collet, 2015; nvc, [n.d.]) can recover data to the exact precision, they suffer from limited compression ratios in floating-point scientific data. On the other hand, lossy compressors for natural vision data and time-series databases, such as JPEGs (Wallace, 1991; Taubman and Marcellin, 2002), H.26x series (Wiegand et al., 2003; Sze et al., 2014; Bross et al., 2021), SummaryStore (Agrawal and Vulimiri, 2017) cannot provide a quantifiable bound on the errors in the decompressed data, leading to uncertainties in the downstream data visualization and analytics. Under these circumstances, error-bounded lossy compression has been proposed as a viable way to reduce the data size while providing guaranteed error control. For instance, ModelarDB (Jensen et al., 2018; Kejser Jensen et al., 2019) is a typical example of an error-bounded compression framework for time-series databases.
@@
The same trend is observed in scientific data management systems. Over the past years, several error-bounded lossy compressors (Liang et al., 2018b, 2023; Lindstrom, 2014; Li et al., 2023; Liu et al., 2024; Tian et al., 2020, 2021) have been developed and adopted by multiple scientific domains, as they can effectively accelerate data management in scientific applications. These compressors feature strict error control in the decompressed data based on user requirements, so domain scientists can control the impact of lossy compression according to their specific needs. However, most compressors only provide error bounds on the raw data, even though the scientists care more about the errors in their derived quantities or downstream analyses. Such derived information, also known as Quantities of Interest (QoIs), is of utmost importance to the fidelity and integrity of scientific discoveries. Typical QoIs include symbolistic functions (Jiao et al., 2022), derivatives and gradients (Su et al., 2022), integrals (Gong et al., 2021a), and topological features (Liang et al., 2022; Yan et al., 2023; Xia et al., 2024). Failure to quantify the distortions in QoIs may lead to misinterpretation of the data and even falsified discoveries, significantly limiting the use of error-bounded lossy compressors in practice.
@@
A few QoI-preserving compressors have already been proposed and applied in scientific applications to bridge the gap. For instance, MGARD (Ainsworth et al., 2019) has enabled the support for bounded linear QoIs, which is further integrated into fusion applications to preserve critical derived quantities such as density (Gong et al., 2021b); variations of SZ have been proposed to tackle the preservation of symbolistic QoIs (Jiao et al., 2022), contour tree (Yan et al., 2023), and critical points in vector fields (Liang et al., 2022; Xia et al., 2024). Despite those efforts, several significant challenges remain unaddressed in the context of QoI reservation. (1) Existing QoI-preserving lossy compressors are finely crafted analytical solutions for particular QoIs, which lose generalization and extensibility to more diverse QoI formats. (2) Most existing methods rely on custom point-wise error bounds to enable the preservation of QoIs, but effectively determining or storing these error bounds is non-trivial. (3) Existing approaches usually tightly couple the QoI-preserving mechanism and compression pipeline, which prohibits using more advanced or newly developed compression methods and limits the efficiency.
To address those challenges, we propose QPET, a novel, versatile, and portable framework to tackle the efficient preservation of symbolistic QoIs. In particular, we notice that various symbolistic QoIs can be modeled in mathematical formats composed of differentiable functions. Moreover, the numerical approximations of the QoIs with a simplified format (e.g., Taylor expansion with second-order derivatives) can provide a unified, efficient, and well-performing algorithmic routine to determine the point-wise error bounds. Thus, we can develop a QoI-preserving framework, which can easily adapt to diverse QoIs and the compression pipelines adopted by lossy compressors in a highly optimized scheme. In summary, our contributions are three-fold:
@@
  • We integrate QPET into 3 state-of-the-art error-bounded lossy compressors and evaluate them with solid benchmarks. Experimental results demonstrate that QPET-integrated compressors yield significant compression ratio (CR) and throughput gains in QoI-preserving compression tasks. Under the same QoI error threshold, QPET-integrated compressors achieve up to 1000% CR improvements over the general-purpose compressors and up to 133% CR improvements over existing QoI-preserving lossy compressors. QPET-integrated compressors can achieve 2x to 10x compression speedups over existing QoI-preserving scientific error-bounded lossy compression solutions.
We organize the rest of this paper as follows: Section 2 discusses the related work of this paper. In Section 3, we formulate and clarify our research motivation and target. In Section 4, we propose the high-level design of QPET, specifying its components, overall algorithm, and integration. Section 5 demonstrates the detailed designs and algorithms in QPET. In Section 6, we present the evaluations of QPET. Section 7 concludes this work and discusses future plans.
@@ 2. Related Work
This section discusses the related work of this paper in three aspects, covering traditional compression techniques used in database systems, emerging lossy compressors for scientific data, and state-of-the-art QoI-preserving lossy compression methods.
@@
Traditional compression techniques for database systems: Various data compressors have been proposed for large-scale databases in diverse data domains and formats. Generic lossless compressors such as GZIP (Deutsch, 1996), ZSTD (Collet, 2015), and LZ4 (Masui et al., 2015) are widely used to compress various types of data to reduce storage requirements (Wiegand et al., 2003; Sze et al., 2014; Pelkonen et al., 2015; Jensen et al., 2018) and accelerate queries (Chen et al., 2001; Arion et al., 2007; Zhang et al., 2018, 2021). With the increasing amount of floating-point data in applications, several compressors have been proposed to deal specifically with floating-point data. A typical example is ndZip (Knorr et al., 2021a, b), which leverages Lorenzo prediction (Ibarria et al., 2003) and residual encoding to compress floating-point data with high efficiency. To handle different types of data, Gorilla (Pelkonen et al., 2015) and Chimp (Liakos et al., 2022) have been designed for time-series data, and Buff (Liu et al., 2021) has been tailored to process low-precision data. Despite their ability to recover the exact data, lossless compressors usually exhibit low compression ratios, which severely limits their use in practice.
@@
To address the limited compression ratios in lossless compressors, lossy compression has been proposed as an alternative way to reduce data in database systems. Regarding lossy database compression methods, ModelarDB (Jensen et al., 2018; Kejser Jensen et al., 2019) summaryStore (Agrawal and Vulimiri, 2017) are designed for time-series data, which leverage data-processing techniques including PMC-mean (Lazaridis and Mehrotra, 2003) and the linear Swing model(Elmeleegy et al., 2009).
@@
Emerging lossy compression techniques for scientific data: To process the high volumes of scientific data while ensuring the quality of decompressed data for accurate post hoc analytics, scientific data compressors require not only decent compression ratios but also strict error control, where error-bounded lossy compression is becoming a promising option. Error-bounded lossy compressors allow users to specify an error bound as a compression parameter and ensure the element-wise difference between the original data and decompressed data is less than the error bound.
@@
SZ3 (Zhao et al., 2021; Liang et al., 2023) is one of the most widely used error-bounded lossy compressors with a prediction-based design. It utilizes dynamic spline functions to approximate the data, followed by a linear-scaling quantization method (Tao et al., 2017) and lossless encoding methods to achieve size reduction with guaranteed error control. Later, HPEZ (Liu et al., 2024) is proposed to significantly improve the quality of SZ3 via more complex interpolation schemes and quality-oriented auto-tuning mechanisms. Another category of error-bounded lossy compressors relies on domain transform for decorrelation. For instance, ZFP (Lindstrom, 2014) is a fast and highly parallel compressor using near-orthogonal block transform, and SPERR (Li et al., 2023) utilizes more complex wavelet transforms to obtain better compression quality at the cost of lower compression throughput. While error-bounded lossy compressors have guaranteed error control on the decompressed data, they cannot provide a quantifiable error bound on the downstream QoIs, as the QoIs could be highly diverse and non-linear.
@@
QoI-preserving lossy compression methods: Variations of error-bounded lossy compressors have been proposed to enable error control for various QoIs. MGARD (Ainsworth et al., 2019) is one of the first compressors to enable QoI preservation, but it can only provide guaranteed error control for bounded-linear QoIs. In (Liang et al., 2020), cpSZ has been proposed to preserve the locations and types of critical points in piece-wise linear vector fields. In particular, it carefully derived sufficient point-wise error bounds based on the target QoI and leveraged it to guide the compression procedure for effective QoI preservation. This idea is further extended to cover the preservation of critical points in multilinear vector fields (Liang et al., 2022), critical points extracted by sign-of-determinant predicates (Xia et al., 2024), contour trees (Yan et al., 2023), and symbolistic QoIs (Jiao et al., 2022).
@@
Despite the rapid development of QoI-preserving lossy compressors, most existing methods suffer from low generalizability and adaptability. In particular, these compressors have to design specific error control mechanisms toward the target QoI, which is hard to generalize to new sets of QoIs. In addition, most existing works are tightly coupled to a specific compression pipeline, making it difficult to integrate with newly developed compression methods. In this work, we propose a portable module that can easily adapt to diverse QoIs and compression methods. This will significantly improve the efficiency and usability of QoI-preserving lossy compression.
@@ 3.1. Quantities of Interest in Scientific Data
As discussed in the previous sections, scientific applications usually require the preservation of QoIs to ensure the integrity of scientific discoveries if lossy compression is used. QoIs represent any derived information that is computed from the raw data, including physical properties (Gong et al., 2021b; Wu et al., 2024), derivatives (Su et al., 2022), and topological features (Yan et al., 2023).
In this work, we mainly target the preservation of symbolistic QoIs , which can be formulated as univariate and multivariate functions of the input data. This covers a wide range of QoIs in scientific applications, including kinetic energy in cosmology, density in fusion energy science, and vector magnitude in climatology.
In the following, we list the set of target QoIs that can be preserved by the proposed framework. They can be broadly categorized into univariate QoIs and multivariate QoIs, as defined below.
@@
Multivariate QoIs: Any differentiable multivariate functions in the format of $F(x_{1},x_{2},...x_{n})$, in which $\{x_{1},x_{2},...x_{n}\}$ are a set of data points, are also supported by QPET. The point set can be a data region or the scalar components of a vector data entry. Typical examples include:
@@
While several existing works managed to preserve certain subsets of the aforementioned QoIs, such as four specific symbolistic QoIs ($x^{2}$, $\log{x}$, $\frac{1}{n}\sum x$, and $\frac{1}{n}\sum x^{2}$(Jiao et al., 2022) and bounded-linear QoIs (Ainsworth et al., 2019), the corresponding methods are based on manual analytical solutions, which cannot be generalized to other QoIs easily because it heavily relies on the analytical form of the QoIs to derive the point-wise error bounds on raw data.
@@ 3.2. QoI-preserving Lossy Compression
Similar to existing works (Jiao et al., 2022; Ainsworth et al., 2019; Liang et al., 2022), we define QoI-preserving lossy compression as lossy compression algorithms that provide guaranteed error control in both the downstream QoIs and the raw data. The mathematical formulation is as follows.
Suppose we have an input data $X=\{x_{i}\}$, and a pair of Compressor and Decompressor $(\operatorname{Cmp},\operatorname{Dec})$ that process $X$ to compressed data $C=\operatorname{Cmp}(X)$ and decompressed data $D=\{d_{i}\}=\operatorname{Dec}(C)$. Given a QoI function $Q$ (which maps data $X$ to another set of data values $Q(X)$), a data error bound $\epsilon$, and a QoI error threshold $\tau$, in QoI-preserving error-bounded lossy compression, we require the errors of data and QoI are both bounded, i.e., $\left\|X-D\right\|_{\infty}\leq\epsilon$ and $\left\|Q(X)-Q(D)\right\|_{\infty}\leq\tau$.
Our research would like to optimize $\operatorname{Cmp}$ and $\operatorname{Dec}$ to maximize the compression ratio. Because most QoIs are non-linear, setting a uniform error bound on each data point will result in divergent QoI errors. To address this issue, existing approaches (Liang et al., 2020, 2022; Jiao et al., 2022) apply different point-wise error bounds on each data point to achieve effective QoI preservation. Nonetheless, these methods focused more on the derivation of the error bounds and overlooked their storage overhead, leading to suboptimal overall compression ratios. To optimize this strategy, selecting the best-fit point-wise error bound (composing a set of $\{\epsilon_{i}\}$) for each data point such that $|x_{i}-x_{i}^{\prime}|\leq\epsilon_{i}$ to maximize compression ratio, we need to jointly optimize the compressed data size by error bounds $\{\epsilon_{i}\}$ ($|\operatorname{Cmp}_{\{\epsilon_{i}\}}(X)|$) and the compressed error-bound size of $\{\epsilon_{i}\}$ themselves ($\left|\operatorname{Cmp}(\{\epsilon_{i}\})\right|$). This optimization problem is shown in Eq. 1 and will be specified in Section 5.

### v4

@@ Abstract.
Error-bounded lossy compression has been widely adopted in many scientific domains because it can address the challenges in storing, transferring, and analyzing unprecedented amounts of scientific data. However, general error-bounded lossy compressors may fail to meet additional quality requirements for downstream analysis, a.k.a. Quantities of Interest (QoIs). This may lead to uncertainties and even misinterpretations in scientific discoveries, significantly limiting the use of lossy compression in practice. In this paper, we propose QPET, a novel, versatile, and portable framework for QoI-preserving error-bounded lossy compression, which overcomes the challenges of modeling diverse QoIs by leveraging numerical strategies. QPET features (1) high portability to multiple existing lossy compressors, (2) versatile preservation to most differentiable univariate and multivariate QoIs, and (3) significant compression improvements in QoI-preservation tasks. Experiments with six real-world datasets demonstrate that integrating QPET into state-of-the-art error-bounded lossy compressors can gain 2x to 10x compression speedups of existing QoI-preserving error-bounded lossy compression solutions, up to 1000% compression ratio improvements to general-purpose compressors, and up to 133% compression ratio improvements to existing QoI-integrated scientific compressors.
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/JLiu-1/QPET-Artifact.
@@ 1. Introduction
Today’s high-performance computing facilities and high-resolution instruments produce vast amounts of scientific data that overwhelm the data storage and transmission systems. According to recent studies, climate simulations generate hundreds of TB of data every 16 seconds (Foster et al., 2017), and fusion applications may produce 1 EB of data in a single run with current- and next-generation exascale computing systems (fus, 2021). In the molecular dynamics (MD) domain, the EXAALT project produces trajectory data with a trillion time steps (exa, 2020), which requires hundreds of terabytes of disk space in a scientific data management system (e.g., HDF5 (hdf, [n.d.])). Those examples pose grand challenges in the underlying data management tasks, including I/O and data transfer, which necessitate effective data compression.
@@
Data compression has been widely adopted in databases and data management systems for reducing data volumes (Wiegand et al., 2003; Sze et al., 2014; Pelkonen et al., 2015; Jensen et al., 2018) and accelerating queries (Chen et al., 2001; Arion et al., 2007; Zhang et al., 2018, 2021). Nonetheless, challenges arise when applying existing compression techniques in the context of scientific data. On the one hand, while lossless compression techniques (Pelkonen et al., 2015; Alakuijala et al., 2018; Lindstrom et al., 2017; Collet, 2015; nvc, [n.d.]) can recover data to the exact precision, they suffer from limited compression ratios in floating-point scientific data. On the other hand, lossy compressors for natural media data and time-series databases, such as JPEGs (Wallace, 1991; Taubman and Marcellin, 2002), H.26x series (Wiegand et al., 2003; Sze et al., 2014; Bross et al., 2021), and SummaryStore (Agrawal and Vulimiri, 2017) cannot provide a quantifiable bound on the decompression errors, leading to uncertainties in the downstream data visualization and analytics. To this end, error-bounded lossy compression has been proposed as a viable way to reduce the data size while providing guaranteed error control. For instance, ModelarDB (Jensen et al., 2018; Kejser Jensen et al., 2019) is a typical example of time-series databases that support error-bounded lossy compression.
@@
The same trend is observed in scientific data management systems. Over the past years, several error-bounded lossy compressors (Liang et al., 2018b, 2023; Lindstrom, 2014; Li et al., 2023; Liu et al., 2024; Tian et al., 2020, 2021) have been developed and adopted by multiple scientific domains, as they can effectively accelerate data management in scientific applications. These compressors feature strict error control based on user requirements, so domain scientists can control the impact of lossy compression according to specific needs. However, most compressors only provide error bounds on the raw data, even though the scientists care more about the errors in their derived quantities or downstream analyses. Such derived information, also known as Quantities of Interest (QoIs), is of utmost importance to the fidelity and integrity of scientific discoveries. Typical QoIs include symbolistic functions (Jiao et al., 2022), derivatives and gradients (Su et al., 2022), integrals (Gong et al., 2021a), and topological features (Liang et al., 2022; Yan et al., 2023; Xia et al., 2024). Failure to quantify the distortions in QoIs may lead to misinterpretation of the data and even falsified discoveries, significantly limiting the use of error-bounded lossy compressors in practice.
@@
A few QoI-preserving compressors have already been proposed and applied in scientific applications to bridge the gap. For instance, MGARD (Ainsworth et al., 2019) has enabled the support for bounded linear QoIs, which is further integrated into fusion applications to preserve critical derived quantities (Gong et al., 2021b); variations of SZ have been proposed to tackle the preservation of symbolistic QoIs (Jiao et al., 2022), contour tree (Yan et al., 2023), and critical points in vector fields (Liang et al., 2022; Xia et al., 2024). Despite those efforts, several significant challenges remain unaddressed in the context of QoI reservation. (1) Existing QoI-preserving lossy compressors are finely crafted for particular QoIs, which lose generalization and extensibility to more diverse QoI formats. (2) Most existing methods rely on custom point-wise error bounds to enable the preservation of QoIs, but effectively determining or storing these error bounds is non-trivial. (3) Existing approaches usually tightly couple the QoI-preserving mechanism and compression pipeline, which fail to leverage newly developed compression methods. To address those challenges, we propose QPET, a novel, versatile, and portable framework for efficiently preserving symbolistic QoIs. In particular, we notice that various symbolistic QoIs can be modeled in mathematical formats composed of differentiable functions. Moreover, the numerical approximations of the QoIs with a simplified format (e.g., Taylor expansion with second-order derivatives) can provide a unified, efficient, and well-performing algorithmic routine to determine the point-wise error bounds. Thus, we develop a highly optimized QoI-preserving framework that can easily adapt to diverse QoIs and scientific lossy compression pipelines. In summary, our contributions are three-fold:
@@
  • We integrate QPET into 3 state-of-the-art error-bounded lossy compressors. Experimental results demonstrate that QPET yields significant compression ratio (CR) and throughput gains in QoI-preserving compression tasks. Under the same QoI error threshold, QPET-integrated compressors achieve up to 1000% CR improvements over the general-purpose compressors and up to 133% CR improvements over existing QoI-preserving lossy compressors. It also achieves 2x to 10x compression speedups over existing QoI-preserving scientific lossy compression solutions.
We organize the rest of this paper as follows: Section 2 discusses the related work. In Section 3, we formulate and clarify our research motivation and target. In Section 4, we propose the high-level design of QPET. Section 5 demonstrates the detailed designs and algorithms in QPET. In Section 6, we present the evaluations of QPET. Section 7 concludes this work and discusses future plans.
@@ 2. Related Work
Traditional compression techniques for database systems: Various data compressors have been proposed for large databases in diverse data domains/formats. Generic lossless compressors such as GZIP (Deutsch, 1996), ZSTD (Collet, 2015), and LZ4 (Masui et al., 2015) are widely used to reduce storage requirements (Wiegand et al., 2003; Sze et al., 2014; Pelkonen et al., 2015; Jensen et al., 2018) and accelerate queries (Chen et al., 2001; Arion et al., 2007; Zhang et al., 2018, 2021). Several lossless compressors have also been proposed to deal specifically with the increasing amount of floating-point data. A typical example is ndZip (Knorr et al., 2021a, b), which leverages Lorenzo prediction (Ibarria et al., 2003) and residual encoding to compress floating-point data with high efficiency. To handle different types of data, Gorilla (Pelkonen et al., 2015) and Chimp (Liakos et al., 2022) have been designed for time-series data, and Buff (Liu et al., 2021) has been tailored to process low-precision data. Despite their ability to recover the exact data, lossless compressors exhibit low compression ratios, severely limiting their use in practice.
@@
To address the limited compression ratios in lossless compressors, lossy compression has been proposed as an alternative way to reduce data in database systems. Regarding lossy database compression methods, ModelarDB (Jensen et al., 2018; Kejser Jensen et al., 2019) and summaryStore (Agrawal and Vulimiri, 2017) are designed for time-series data, which leverage data-processing techniques including PMC-mean (Lazaridis and Mehrotra, 2003) and the linear Swing model(Elmeleegy et al., 2009).
@@
Emerging lossy compression techniques for scientific data: To process the high volumes of scientific data while preserving the data quality for accurate post hoc analytics, scientific data compressors require not only decent compression ratios but also strict error control, where error-bounded lossy compression is becoming a promising option. It allows users to specify an error bound as a parameter and ensures the element-wise difference between the original data and decompressed data is less than the error bound.
@@
SZ3 (Zhao et al., 2021; Liang et al., 2023) is one of the most widely used error-bounded lossy compressors with a prediction-based design. It utilizes dynamic spline interpolations to approximate the data, followed by linear-scaling quantization  (Tao et al., 2017) and lossless encoding to reduce data size with guaranteed error control. Later, QoZ/HPEZ (Liu et al., 2022, 2024) is proposed to significantly improve the quality of SZ3 via more complex interpolation schemes and quality-oriented auto-tuning mechanisms. Another category of error-bounded lossy compressors relies on domain transform for decorrelation. For instance, ZFP (Lindstrom, 2014) is a fast compressor using near-orthogonal block transform, and SPERR (Li et al., 2023) utilizes more complex wavelet transforms to obtain better compression quality at the cost of lower compression throughput. While error-bounded lossy compressors have guaranteed data error control, they cannot provide a quantifiable error bound on the downstream QoIs, as the QoIs could be highly diverse and non-linear.
@@
QoI-preserving lossy compression methods: Several solutions have been proposed to enable error control for various QoIs. MGARD (Ainsworth et al., 2019) is one of the first compressors to enable QoI preservation, but it can only provide guaranteed error control for bounded-linear QoIs. In (Liang et al., 2020), cpSZ has been proposed to preserve the locations and types of critical points in piece-wise linear vector fields. In particular, it carefully derived sufficient point-wise error bounds based on the target QoI and leveraged them to guide the compression procedure for effective QoI preservation. This idea is further extended to cover the preservation of critical points in multilinear vector fields (Liang et al., 2022), critical points extracted by sign-of-determinant predicates (Xia et al., 2024), contour trees (Yan et al., 2023), and symbolistic QoIs (Jiao et al., 2022).
@@
Most existing QoI-preserving lossy compressors suffer from low generalizability and adaptability. In particular, they feature specific error control mechanisms toward the target QoI, which is hard to generalize to new sets of QoIs. In addition, most existing works are tightly coupled to a specific compression pipeline, making it difficult to integrate with newly developed compression methods. In this work, we propose a portable module that can easily adapt to diverse QoIs and compression methods. This will significantly improve the efficiency and usability of QoI-preserving lossy compression.
@@ 3.1. Quantities of Interest in Scientific Data
As previously discussed, scientific applications usually require the preservation of QoIs to ensure the integrity of scientific discoveries if lossy compression is used. QoIs represent any derived information that is computed from the raw data, including physical properties (Gong et al., 2021b; Wu et al., 2024), derivatives (Su et al., 2022), and topological features (Yan et al., 2023).
In this work, we target the preservation of symbolistic QoIs , which can be formulated as mathematical functions of the input data. This covers various QoIs in scientific domains, including kinetic energy in cosmology, density in fusion energy science, and vector magnitude in climatology. We list the categorizations of target QoIs that the proposed framework can preserve as follows:
@@
Multivariate QoIs: Any differentiable multivariate functions in the format of $F(x_{1},x_{2},...x_{n})$, in which $\{x_{1},x_{2},...x_{n}\}$ are a set of data points, are also supported by QPET. The point set can be a data region or a vector data entry. Typical examples include:
@@
While several existing works managed to preserve certain subsets of the aforementioned QoIs (Jiao et al., 2022; Ainsworth et al., 2019), they cannot be generalized to other QoIs easily because they heavily rely on the analytical form of the QoIs to make analysis.
@@ 3.2. QoI-preserving Lossy Compression
We mathematically define QoI-preserving lossy compression as providing guaranteed error control in both the downstream QoIs and the raw data, which is similar to existing works (Jiao et al., 2022; Ainsworth et al., 2019; Liang et al., 2022).
Given an input data $X=\{x_{i}\}$, and a pair of compressor and decompressor $(\operatorname{Cmp},\operatorname{Dec})$ that process $X$ to compressed data $C=\operatorname{Cmp}(X)$ and decompressed data $D=\{d_{i}\}=\operatorname{Dec}(C)$, for a QoI function $Q$ (which maps $X$ to another set of values $Q(X)$), a data error bound $\epsilon$, and a QoI error threshold $\tau$, in QoI-preserving error-bounded lossy compression, we require the errors of data and QoI are both bounded, i.e., $\left\|X-D\right\|_{\infty}\leq\epsilon$ and $\left\|Q(X)-Q(D)\right\|_{\infty}\leq\tau$.
Our research would like to optimize $\operatorname{Cmp}$ and $\operatorname{Dec}$ to maximize the compression ratio. Because most QoIs are non-linear, setting a uniform error bound on each data point will result in divergent QoI errors. To address this issue, existing approaches (Liang et al., 2020, 2022; Jiao et al., 2022) apply different point-wise error bounds on each data point to achieve effective QoI preservation. Nonetheless, these methods focused more on deriving the error bounds and overlooked their storage overhead, leading to suboptimal overall compression ratios. To optimize this strategy, selecting the best-fit point-wise error bound (composing a set of $\{\epsilon_{i}\}$) for each data point such that $|x_{i}-x_{i}^{\prime}|\leq\epsilon_{i}$ to maximize compression ratio, we need to jointly optimize the compressed data size ($|\operatorname{Cmp}_{\{\epsilon_{i}\}}(X)|$) and the compressed error-bound size of $\{\epsilon_{i}\}$ themselves ($\left|\operatorname{Cmp}(\{\epsilon_{i}\})\right|$). This optimization problem is shown in Eq. 1 and will be specified in Section 5.